# Needle in the Haystack for Memory Based Large Language Models

### v1

@@ Abstract
In this paper, we demonstrate the benefits of using memory augmented Large Language Model (LLM) architecture in improving the recall abilities of facts from a potentially long context. As a case study we test larimar, a recently proposed LLM architecture which augments a LLM decoder with an external associative memory, on several long-context recall tasks, including passkey and needle-in-the-haystack tests. We demonstrate that the external memory can be adapted at test time to handle contexts much longer than those seen during training, while keeping readouts from the memory recognizable to the trained decoder and without increasing GPU memory footprint. Compared to alternative architectures for long-context recall tasks with models of a comparable parameter count, larimar is able to maintain strong performance without any task-specific training.
@@
Subhajit Chaudhury1, Soham Dan1, Payel Das1, Georgios Kollias1, Elliot Nelson1 1IBM Research Correspondence: enelson@ibm.com

### v2

@@ Abstract
Current large language models (LLMs) often perform poorly on simple fact retrieval tasks. Here we investigate if coupling a dynamically adaptable external memory to a LLM can alleviate this problem. For this purpose, we test Larimar, a recently proposed language model architecture which uses an external associative memory, on long-context recall tasks including passkey and needle-in-the-haystack tests. We demonstrate that the external memory of Larimar, which allows fast write and read of an episode of text samples, can be used at test time to handle contexts much longer than those seen during training. We further show that the latent readouts from the memory (to which long contexts are written) control the decoder towards generating correct outputs, with the memory stored off of the GPU. Compared to existing transformer-based LLM architectures for long-context recall tasks that use larger parameter counts or modified attention mechanisms, a relatively smaller size Larimar is able to maintain strong performance without any task-specific training or training on longer contexts.
@@
Elliot Nelson1, Georgios Kollias2, Payel Das3, Subhajit Chaudhury4, Soham Dan5 IBM Research
11footnotetext: enelson@ibm.com22footnotetext: gkollias@us.ibm.com33footnotetext: daspa@us.ibm.com44footnotetext: subhajit@ibm.com55footnotetext: soham.dan@ibm.com